OpenStudy (anonymous):

The cost of £C, of hiring a lorry to travel n miles is given by C=a+bn where a and b are consonants The cost of travelling 100 miles is £82. The cost of travelling 250 miles is £157. a) Find the values of a and b b)Work out the cost of hiring a lorry to travel 300 miles. Please can you show the working out as well.

OpenStudy (anonymous):

Solve the following for a and b:\[\{82=a+b *100,157=a+b* 250\} \]\[\left\{a=32,b=\frac{1}{2}\right\} \]

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

Ur welcome. Tank you for the answer. Great help!!! :)